Determine the number of CH2Cl2 molecules in 25.4 g CH2Cl2.

Determine the number of  molecules in 25.4  .

2.00×10^25 molecules
0.299 molecules
1.80×10^23 molecules
1.62×10^27 molecules
Answer #1

1 mol of any substance contains 6.022*1023 molecules/atoms. (Avogadro law)

Calculating the moles of given CH2cl2

Number of moles = given weight(g)/molecular weight (gmol-1)

= 25.4g/85gmol-1. = 0.2988 mol

So number of molecules in 0.2988 mol

= 0.2988*6.022*1023 molecules = 1.79*1023 molecules

~1.80*1023 molecules